OpenStudy (anonymous):

Evaluate v^2-3 when v=5

OpenStudy (anonymous):

I came up with 22

OpenStudy (anonymous):

25-3 = 22 correct

OpenStudy (anonymous):

kool

OpenStudy (anonymous):

just 22

OpenStudy (anonymous):

yes 5^2 = 25 25-3 = 22